If Chuck Norris was riding a Razor Scooter down Gemini Blvd., and his time at 202.6 meters was 30 seconds and his time at 245 meters was 42 seconds, then what was his average velocity?
Answer #1

Average velocity= Displacement/Time interval

v=average velocity

s= Displacemen = (Final position - initial position)= (245-202.6)= 42.4 m

t= Time interval= (42-30)= 12 sec

v=s/t= 42.4/12= 3.53

Average velocity= 3.53 m/sec